# Vizing’s Theorem in Near-Linear Time

### v1

@@ 2.2 Vizing fans and chains
Vizing Fans: We now define the notion of Vizing fans, which has been used extensively in the edge coloring literature [Viz64, GNK+85, Sin19].
@@
We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{0},c_{0}),\dots,(v_{k},c_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$. We refer to $c_{i}$ as the color of $v_{i}$ within $\operatorname*{\mathbf{F}}$. A crucial property is that we can rotate colors around the Vizing fan $\operatorname*{\mathbf{F}}$ by setting $\chi(u,v_{1})\leftarrow c_{1},\dots,\chi(u,v_{i-1})\leftarrow c_{i-1},\chi(u,v_{i})\leftarrow\bot$ for any $i\in[k]$. We say that $\operatorname*{\mathbf{F}}$ is a trivial Vizing fan if $c_{k}\in\mathsf{miss}_{\chi}(u)$. Note that, if $\operatorname*{\mathbf{F}}$ is trivial, we can immediately extend the coloring $\chi$ to $(u,v_{1})$ by rotating all the colors around $\operatorname*{\mathbf{F}}$ and setting $\chi(u,v_{k})\leftarrow c_{k}$.
@@
1 For each $x\in V$, let $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)$
@@ 7       $k\leftarrow k+1$
1 if $\operatorname*{\mathbf{F}}$ is trivial then
@@ 2       $\chi(u,v_{i})\leftarrow c_{i}$ for all $i\in[k]$
[…]
@@ 3.1 U-Fans and U-Edges
We say that $u$ is the center of $\operatorname*{\boldsymbol{e}}$ and that $\alpha$ is the active color of $\operatorname*{\boldsymbol{e}}$. For notational convenience, we also say that the u-edge $\operatorname*{\boldsymbol{e}}$ is $\alpha$-primed and define $c_{\operatorname*{\boldsymbol{e}}}(u):=\alpha$ and $c_{\operatorname*{\boldsymbol{e}}}(v)=\bot$. [^3] Whenever we refer to an ‘uncolored edge $e$’, we are referring to an edge $e\in E$ such that $\chi(e)=\bot$, whereas a ‘u-edge $\operatorname*{\boldsymbol{e}}$’ always refers to the object from LABEL:{def:u-edge} and is denoted in bold.
@@ 3.2 Data Structures
In Section 7, we describe the data structures that we use to implement our algorithm. On top of the standard data structures used to maintain the edge coloring $\chi$, which allows us to implement Algorithms 1 and 2 efficiently, we also use data structures that allow us to efficiently maintain a separable collection $\operatorname*{\mathcal{U}}$. More specifically, the data structures that we use to implement a separable collection $\operatorname*{\mathcal{U}}$ support the following queries.
@@
[…]
@@
A Remark on Randomization: In order to get good space complexity and running time simultaneously, we implement the standard data structures (used for Algorithms 1 and 2) and the data structure supporting the preceding queries using hashmaps (see Section 7). The following proposition describes the hashmaps used by our data structures. [^5] We could have even used the construction of [Kus22] to obtain exponentially small error probability in this case, but since we do not need this stronger guarantee we stick with the simpler work of [DadH90].
@@ 3.3 Vizing Fans in Separable Collections
Let $\operatorname*{\mathcal{U}}$ be a separable collection and $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ be a Vizing fan. We say that the Vizing fan $\operatorname*{\mathbf{F}}$ if $\operatorname*{\mathcal{U}}$-avoiding if $c_{i}\in\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each leaf $v_{i}\in\operatorname*{\mathbf{F}}$.
@@
By 3.5, we can always find a collection of colors $\{\mathsf{clr}(x)\}_{x\in V}$ such that $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ for each $x\in V$. If we construct a Vizing fan $\operatorname*{\mathbf{F}}$ by calling Vizing-Fan and using such a collection of colors within the algorithm, then it follows that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding. By combining the standard implementation of Algorithm 1 with the queries described in Section 3.2 we can do this in $O(\Delta)$ time. ∎
@@ 4 The Main Algorithm
Our final algorithm consists of two main components. The first component is an algorithm called Construct-U-Fans that takes a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda$ uncolored edges and either extends $\chi$ to $\Omega(\lambda)$ of these edges or modifies the coloring to construct a separable collection of $\Omega(\lambda)$ u-fans. Lemma 4.1 summarizes the behaviour of this algorithm.
@@
Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, the algorithm Construct-U-Fans does one of the following in $O((m+\Delta\lambda)\log n)$ time:
@@
The second component is an algorithm called Color-U-Fans that takes a collection of $\lambda$ u-fans and extends the coloring to $\Omega(\lambda)$ of the edges in the u-fans. Lemma 4.2 summarizes the behaviour of this algorithm.
@@
In Sections 5 and 6, we describe these algorithms and prove Lemmas 4.1 and 4.2 respectively. Using these lemmas, we now show how to efficiently extend an edge coloring $\chi$ to the remaining uncolored edges.
@@
Let $U$ denote the set of edges that are uncolored by $\chi$. We can then apply Lemma 4.1 to either extend $\chi$ to a constant fraction of the edges in $U$ or construct a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$ in $O((m+\Delta\lambda)\log n)$ time. In the second case, we can then apply Lemma 4.2 to color $\Omega(\lambda)$ of the edges contained in these u-fans in $O(m\log n)$ time w.h.p. Thus, we can extend $\chi$ to a constant fraction of the uncolored edges in $O((m+\Delta\lambda)\log n)$ time w.h.p. After repeating this process $O(\log\lambda)\leq O(\log n)$ many times, no edges remain uncolored. Thus, we can extend the coloring $\chi$ to the entire graph in $O((m+\Delta\lambda)\log^{2}n)$ time w.h.p. ∎
@@
5 The Algorithm Construct-U-Fans (Proof of Lemma 4.1)
@@
1 Construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$
@@ 3       $\mathcal{F}\leftarrow\textnormal{{Prune-Vizing-Fans}}(\operatorname*{\mathcal{U}},\alpha)$
[…]
@@ 5.1 The Subroutine Prune-Vizing-Fans
Algorithm Description: The subroutine Prune-Vizing-Fans scans through the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and constructs $\operatorname*{\mathcal{U}}$-avoiding Vizing fans at these u-edges (see Lemma 3.8), maintaining a vertex-disjoint subset $\mathcal{F}$ of these Vizing fans throughout the process. After constructing a Vizing fan $\operatorname*{\mathbf{F}}$ for a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, the subroutine checks if $F$ intersects some other fan in $\mathcal{F}$. If not, it adds $\operatorname*{\mathbf{F}}$ to $\mathcal{F}$. Otherwise, it uses these intersecting Vizing fans to either extend the coloring $\chi$ or construct a u-fan, removing the corresponding Vizing fans and u-edges from $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ respectively.
@@
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the set $\mathcal{F}$ consists of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for the u-edges in $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i}\}\cap\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. [^8] Since we remove edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ throughout this process, $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i-1}\}$ might not be contained in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. This is true trivially for $i=1$. Now, suppose that this is true for $i-1$. If the $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ does not intersect any of the Vizing fans in $\mathcal{F}$, then we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$ and are done.
@@
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the collection $\operatorname*{\mathcal{U}}$ is separable. This is true trivially for $i=1$, since the subroutine does not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning this u-edge. Now, suppose that this is true for $i-1$. It follows from the inductive argument in Lemma 5.1 that the Vizing fans currently in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding. If we do not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning $\operatorname*{\boldsymbol{e}}_{i}$, then we are done. Thus, we assume that this is not the case. The subroutine can modify $\chi$ and $\operatorname*{\mathcal{U}}$ in the following ways:
@@ 5.2 The Subroutine Reduce-U-Edges
[…]
@@
[…]
@@
1 $P_{\operatorname*{\mathbf{F}}}\leftarrow\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq|P_{\operatorname*{\mathbf{F}}}|+1}$
2 Let $(x,y)$ be the $(L+1)^{th}$ edge in $P_{\operatorname*{\mathbf{F}}}$ and $y$ be an endpoint of $P_{\operatorname*{\mathbf{F}}}$

### v3

@@ 4 Preliminaries: Vizing Fans and Vizing Chains
We now define the notion of Vizing fans, which has been used extensively in the edge coloring literature [Viz64, GNK+85, Sin19].
@@
We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$. We refer to $c_{i}$ as the color of $v_{i}$ within $\operatorname*{\mathbf{F}}$. A crucial property is that we can rotate colors around the Vizing fan $\operatorname*{\mathbf{F}}$ by setting $\chi(u,v_{1})\leftarrow c_{1},\dots,\chi(u,v_{i-1})\leftarrow c_{i-1},\chi(u,v_{i})\leftarrow\bot$ for any $i\in[k]$. We say that $\operatorname*{\mathbf{F}}$ is a trivial Vizing fan if $c_{k}\in\mathsf{miss}_{\chi}(u)$. Note that, if $\operatorname*{\mathbf{F}}$ is trivial, we can immediately extend the coloring $\chi$ to $(u,v_{1})$ by rotating all the colors around $\operatorname*{\mathbf{F}}$ and setting $\chi(u,v_{k})\leftarrow c_{k}$.
@@
1For each $x\in V$, let $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)$
@@ 7 $k\leftarrow k+1$
1if $\operatorname*{\mathbf{F}}$ is trivial then
@@ 2 $\chi(u,v_{i})\leftarrow c_{i}$ for all $i\in[k]$
Given a Vizing fan $\operatorname*{\mathbf{F}}$, we denote the path $P$ considered by Algorithm 3 by $\texttt{Vizing-Path}(\operatorname*{\mathbf{F}})$. If the Vizing fan $\operatorname*{\mathbf{F}}$ is trivial, then $\texttt{Vizing-Path}(\operatorname*{\mathbf{F}})$ denotes an empty path $\varnothing$.
@@ 5.1 U-Fans and U-Edges
We say that $u$ is the center of $\operatorname*{\boldsymbol{e}}$ and that $\alpha$ is the active color of $\operatorname*{\boldsymbol{e}}$. For notational convenience, we also say that the u-edge $\operatorname*{\boldsymbol{e}}$ is $\alpha$-primed and define $c_{\operatorname*{\boldsymbol{e}}}(u):=\alpha$ and $c_{\operatorname*{\boldsymbol{e}}}(v)=\bot$. [^4] Whenever we refer to an “uncolored edge ee”, we are referring to an edge $e\in E$ such that $\chi(e)=\bot$, whereas a ‘u-edge $\operatorname*{\boldsymbol{e}}$’ always refers to the object from LABEL:{def:u-edge} and is denoted in bold.
@@
We remark that the second property of this definition is rather important because we need to ensure that different u-components are not interfering with each other when they share common vertices. Check Figure 4 for an illustration.
[…]
Figure 4: This picture shows two u-fans $(u_{1},x_{1},x,*,\beta_{1})$ and $(u_{2},x_{2},x,*,\beta_{2})$ sharing a common vertex $x$. The separable condition requires that $\beta_{1}\neq\beta_{2}$; for instance $\beta_{1},\beta_{2}$ could be magenta and cyan as shown here.
@@ 5.2 Data Structures
In Section 9, we describe the data structures that we use to implement our algorithm. On top of the standard data structures used to maintain the edge coloring $\chi$, which allows us to implement Algorithms 2 and 3 efficiently, we also use data structures that allow us to efficiently maintain a separable collection $\operatorname*{\mathcal{U}}$. More specifically, the data structures that we use to implement a separable collection $\operatorname*{\mathcal{U}}$ support the following queries.
@@
  • $\textsc{Find-Component}_{\operatorname*{\mathcal{U}}}(x,c)$: The input to this query is a vertex $x\in V$ and a color $c\in[\Delta+1]$. In response, the data structure returns the u-component $g\in\operatorname*{\mathcal{U}}$ with $c_{\operatorname*{\boldsymbol{g}}}(x)=c$ if such a u-component exists and outputs fail otherwise.
  • $\textsc{Missing-Color}_{\operatorname*{\mathcal{U}}}(x)$: The input to this query is a vertex $x\in V$. In response, the data structure returns an arbitrary color from the set $\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$. [^5] Note that, since $|C_{\operatorname*{\mathcal{U}}}(x)|<|\mathsf{miss}_{\chi}(x)|$, such a color always exists.
@@
A Remark on Randomization: In order to get good space complexity and running time simultaneously, we implement the standard data structures (used for Algorithms 2 and 3) and the data structure supporting the preceding queries using hashmaps (see Section 9). The following proposition describes the hashmaps used by our data structures. [^6] We could have even used the construction of [Kus22] to obtain exponentially small error probability in this case, but since we do not need this stronger guarantee we stick with the simpler work of [DadH90].
@@ 5.3 Vizing Fans in Separable Collections
Let $\operatorname*{\mathcal{U}}$ be a separable collection and $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ be a Vizing fan. We say that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding if $c_{i}\in\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each leaf $v_{i}\in\operatorname*{\mathbf{F}}$.
@@
By 5.5, we can always find a collection of colors $\{\mathsf{clr}(x)\}_{x\in V}$ such that $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ for each $x\in V$. If we construct a Vizing fan $\operatorname*{\mathbf{F}}$ by calling Vizing-Fan and using such a collection of colors within the algorithm, then it follows that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding. By combining the standard implementation of Algorithm 2 with the queries described in Section 5.2 we can do this in $O(\Delta)$ time. ∎
@@ 6 The Main Algorithm
We are now ready to present our main technical result, which is a slightly weaker version of Theorem 1.1, and focuses on achieving a near-linear time algorithm for $(\Delta+1)$ edge coloring (instead of the exact $O(m\log{n})$ time in Theorem 1.1; see the remark after that theorem). We will then use this theorem in Section 10 to conclude the proof of Theorem 1.1.
Theorem 6.1.
Our main algorithm consists of two main components. The first component is an algorithm called Construct-U-Fans that takes a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda$ uncolored edges and either extends $\chi$ to $\Omega(\lambda)$ of these edges or modifies the coloring to construct a separable collection of $\Omega(\lambda)$ u-fans. Lemma 6.2 summarizes the behavior of this algorithm.
@@
Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, the algorithm Construct-U-Fans does one of the following in $O(m+\Delta\lambda)$ time:
@@
The second component is an algorithm called Color-U-Fans that takes a collection of $\lambda$ u-fans and extends the coloring to $\Omega(\lambda)$ of the edges in the u-fans. Lemma 6.3 summarizes the behavior of this algorithm. The reader may find it helpful to keep in mind that the algorithm for Lemma 6.3 is a generalization of algorithm for Lemma 3.2 in Section 3.
@@
In Sections 7 and 8, we prove Lemmas 6.2 and 6.3 respectively. Using these lemmas, we now show how to efficiently extend an edge coloring $\chi$ to the remaining uncolored edges.
@@
Let $U$ denote the set of edges that are uncolored by $\chi$. We can then apply Lemma 6.2 to either extend $\chi$ to a constant fraction of the edges in $U$ or construct a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$ in $O(m+\Delta\lambda)$ time. In the second case, we can then apply Lemma 6.3 to color $\Omega(\lambda)$ of the edges contained in these u-fans in $O(m\log n)$ time w.h.p. Thus, we can extend $\chi$ to a constant fraction of the uncolored edges in $O(m\log n+\Delta\lambda)$ time w.h.p. After repeating this process $O(\log\lambda)\leq O(\log n)$ many times, no edges remain uncolored. Thus, we can extend the coloring $\chi$ to the entire graph in $O((m+\Delta\lambda)\log^{2}n)$ time w.h.p. ∎
@@
7 The Algorithm Construct-U-Fans: Proof of Lemma 6.2
@@
1Construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$
@@ 3 $\mathcal{F}\leftarrow\textnormal{{Prune-Vizing-Fans}}(\operatorname*{\mathcal{U}},\alpha)$
4 $\textnormal{{Reduce-U-Edges}}(\operatorname*{\mathcal{U}},\alpha,\mathcal{F})$
@@ 7.1 The Subroutine Prune-Vizing-Fans
Algorithm Description: The subroutine Prune-Vizing-Fans scans through the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and constructs $\operatorname*{\mathcal{U}}$-avoiding Vizing fans at these u-edges (see Lemma 5.8), maintaining a vertex-disjoint subset $\mathcal{F}$ of these Vizing fans throughout the process. After constructing a Vizing fan $\operatorname*{\mathbf{F}}$ for a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, the subroutine checks if $\operatorname*{\mathbf{F}}$ intersects some other fan in $\mathcal{F}$. If not, it adds $\operatorname*{\mathbf{F}}$ to $\mathcal{F}$. Otherwise, it uses these intersecting Vizing fans to either extend the coloring $\chi$ or construct a u-fan, removing the corresponding Vizing fans and u-edges from $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ respectively. See Figure 5 for an illustration.
[…]
Figure 5: In this picture, we look at the Vizing fan around an uncolored edge $(u_{1},v_{1})$, and it intersects with the Vizing fan of another edge $(u_{2},v_{2})$ currently residing in $\mathcal{F}$ at vertex $w$. Then, we can rotate both Vizing fans and pair these two uncolored edges as a u-fan.
@@
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the set $\mathcal{F}$ consists of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for the u-edges in $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i}\}\cap\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. [^9] Since we remove edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ throughout this process, $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i-1}\}$ might not be contained in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. This is true trivially for $i=1$. Now, suppose that this is true for i1i-1. If the $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ does not intersect any of the Vizing fans in $\mathcal{F}$, then we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$ and are done.
@@
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the collection $\operatorname*{\mathcal{U}}$ is separable. This is true trivially for $i=1$, since the subroutine does not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning this u-edge. Now, suppose that this is true for i1i-1. It follows from the inductive argument in Lemma 7.1 that the Vizing fans currently in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding. If we do not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning $\operatorname*{\boldsymbol{e}}_{i}$, then we are done. Thus, we assume that this is not the case. The subroutine can modify $\chi$ and $\operatorname*{\mathcal{U}}$ in the following ways:
@@ 7.2 The Subroutine Reduce-U-Edges
Algorithm Description: The subroutine Reduce-U-Edges begins by activating all of the trivial Vizing fans in $\mathcal{F}$ and then removing them from $\mathcal{F}$. It then constructs Vizing chains at each of the remaining Vizing fans in $\mathcal{F}$ and explores the alternating paths $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$ ‘in parallel’ (meaning, one edge at a time across all these paths). The subroutine continues this process until it either (1) reaches the end of one of these paths, or (2) two of these paths intersect. In the first case, it identifies the Vizing fan $\operatorname*{\mathbf{F}}\in\mathcal{F}$ whose Vizing chain is fully explored and calls $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ before removing $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$. In the second case, it identifies the Vizing fans $\operatorname*{\mathbf{F}},\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ corresponding to the intersecting Vizing chains and either extends the coloring $\chi$ to one more uncolored edge or creates a u-fan by shifting uncolored edges down these Vizing chains, before removing $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$. It repeats this process until at least half of the Vizing fans have been removed from $\mathcal{F}$.
@@
    For each $\operatorname*{\mathbf{F}}\in\mathcal{F}$, $P_{\operatorname*{\mathbf{F}}}$ is the length $|P_{\operatorname*{\mathbf{F}}}|$ prefix of the alternating path $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$, which we denote by $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq|P_{\operatorname*{\mathbf{F}}}|}$.
@@
Immediately after starting a round, the subroutine increases the value of $L$ by $1$, invalidating 7.6. To restore this invariant, we need to update each of the prefix paths $P_{\operatorname*{\mathbf{F}}}$, which in turn may require us to update the other objects maintained by the subroutine to ensure that the other conditions of 7.5 are satisfied. The subroutine does this by scanning through each $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and calling $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$, which we describe formally in Algorithm 5. The subroutine performs these rounds until $|\mathcal{F}|\leq\lambda_{\alpha}/2$, at which point it terminates.
@@
2 Let $(x,y)$ be the $L^{th}$ edge in $P_{\operatorname*{\mathbf{F}}}$ and $y$ be an endpoint of $P_{\operatorname*{\mathbf{F}}}$